What is the concentration of an Ammonia (NH3) solution when 25.0 g of Ammonia are dissolved in enough water to produce 650.0 mL of solution?

A. 2.26 M
B. 0.719 moles
C. 1.47 M
D. 1.47 moles
E. None of the Above
1
Expert's answer
2018-03-22T08:41:39-0400
Solution
n = 25 g / 17.031 g/mol
C = n / V = 25 g / 17.031 g/mol / 0.650 L = 2.258 M ≈ 2.26 M
Answer: A